### v1

# Geometry-Aware Edge Pooling for Graph Neural Networks

Abstract
Graph Neural Networks (GNNs) have shown significant success for graph-based tasks. Motivated by the prevalence of large datasets in real-world applications, pooling layers are crucial components of GNNs. By reducing the size of input graphs, pooling enables faster training and potentially better generalisation. However, existing pooling operations often optimise for the learning task at the expense of fundamental graph structures and interpretability. This leads to unreliable performance across varying dataset types, downstream tasks and pooling ratios. Addressing these concerns, we propose novel graph pooling layers for structure-aware pooling via edge collapses. Our methods leverage diffusion geometry and iteratively reduce a graph’s size while preserving both its metric structure and structural diversity. We guide pooling using magnitude, an isometry-invariant diversity measure, which permits us to control the fidelity of the pooling process. Further, we use the spread of a metric space as a faster and more stable alternative ensuring computational efficiency. Empirical results demonstrate that our methods (i) achieve superior performance compared to alternative pooling layers across a range of diverse graph classification tasks, (ii) preserve key spectral properties of the input graphs, and (iii) retain high accuracy across varying pooling ratios.
1 Introduction
Graph pooling layers are important components of successful GNN architectures. They are implemented alongside convolutional layers to reduce the size of graph representations during training. Pooling thus enables GNNs to scale to large and complex real-world graphs. However, the choice of pooling method strongly influences downstream-applications and task-performance. In fact, the question of which graph properties to preserve during pooling [44, 32, 34], just as the question on the nature and quality of graph datasets in graph learning [13], remains an ongoing debate. It is thus crucial to design expressive, efficient, and interpretable pooling layers that are capable of reliably encoding task-relevant information while reducing the size of input graphs. Most graph pooling literature takes a node-centric view [32]. However, this focus on node-centric rather than edge-centric pooling often leads to the loss of important structural information. Common pooling methods either drop nodes or optimise for a node clustering while treating graph connectivities as a secondary objective. As visualised in Figure 1 and further explored in our work, this frequently leads to counter-intuitive pooling decisions that fail to retain key geometric structures in the graph.
Addressing these concerns, topological and geometric descriptors of graphs are uniquely poised to interoperate structural information into graph pooling. Throughout this work, we treat graphs as metric spaces and assess their geometry via diffusion distances, which naturally work alongside message passing to effectively encode key graph structures. Motivated by ongoing research on novel geometric invariants, we find that generalised measures of size and diversity are especially promising candidates for guiding graph pooling. In particular, we use the magnitude of a metric space, which measures a graph’s structural diversity, to control the loss of structural information during edge pooling [30]. Our work is motivated by successful applications of magnitude across a range of machine learning tasks, such as the evaluation of diversity [31] for generative modelling, boundary detection for images [2], and the study of generalisation for neural networks [3, 4]. Building up on this research, we are the first to propose the use of magnitude in the context of graph learning. We further advance on existing applications by investigating an alternative and closely-related measure, known as the spread of a metric space, to greatly improve the computational efficiency of our methods.
Our main contributions are as follows:
  • We propose MagEdgePool and SpreadEdgePool, novel edge-contraction based pooling layers, which preserve graphs’ structural diversity.
  • We propose the spread of a metric space as a faster and more stable alternative to magnitude and ensure that our algorithms can be computed efficiently.
  • We evaluate our methods’ capability to preserve key structural properties during pooling.
  • We demonstrate that our pooling methods consistently perform well at graph classification, surpassing alternative pooling layers across a wide range of datasets and pooling ratios.
2 Background
2.1 Graph Pooling for Graph Neural Networks
Figure 1: Examples of pooled graphs compared across different pooling layers. Our proposed methods, MagEdgePool and SpreadEdgePool, respect the original graphs’ geometry during pooling. Alternative approaches tend to obscure adjacency relationships to varying extents by creating counter-intuitive edges (Graclus, NMF), disconnecting entire portions of the graphs (TopK, SAGPool), or returning dense representations that do not preserve any geometric structure (DiffPool, MinCut).
As an ongoing field of interest for graph learning, a wide range of pooling methods has been proposed. The most successful of these methods often use a hierarchical approach to sequentially coarsen the graph representation while reducing its size [32]. Research has traditionally focused on node-based pooling via node clustering or node dropping. Amongst node drop pooling methods, Node Decimal Pooling (NDP) [9] is non-trainable, while TopK [21, 11] and SAGPool [27], are trainable approaches. However, all these methods inevitably lose information because they do not select all vertices but remove entire sections of the graph during pooling [32]. Node clustering approaches are similarly either non-trainable, such as Graclus [15] and Non-Negative Matrix Factorization (NMF) [6], or trainable, such as MinCUT [8] and DiffPool [45], which output dense graph representations. In comparison to node drop pooling, clustering-based methods usually require high memory costs [32]. Overall, trainable methods have the potential to better optimise for a specific objective, but might overfit to the task at hand, especially for smaller datasets. By contrast, non-trainable methods can act as a stronger inductive bias on the underlying graph representation, and do not introduce additional trainable parameters or optimisation objectives into the GNN [25]. Edge-based pooling methods have been studied less extensively than node-centric pooling [16, 26] despite their promising potential for encoding connectivities in a more faithful and natural manner. EdgePool [16], the most successful edge-based method, uses iterative edge-contraction based on edge scores, which are learned from features of adjacent nodes. However, EdgePool always pools graphs to half their size, making it less flexible than more adaptive methods. Further, it does not explicitly consider a graph’s topology, beyond posing a constraint on not contracting adjacent edges. Thus, selecting edges based on node features can lead to counter-intuitive decisions during pooling, for example by retaining local structures in strongly-connected communities even at high pooling ratios [16].
Interpretability remains a leading concern [32], and non-trainable approaches are capable of being well-performing and traceable baseline pooling methods [25]. We therefore find that there is strong potential for designing geometry-aware edge pooling operations that make interpretable decisions on which aspects of the graph to retain. Figure 1 further illustrates how many of the aforementioned pooling methods inherently fail to preserve key graph structures even for simple toy examples. Trainable methods in this overview were optimised for a spectral loss following Grattarola et al. [25], but this does not ensure interpretable preservation of graphs’ underlying geometry. Addressing these shortcomings, it is of interest to leverage alternative tools from computational geometry, which allow us to quantify the qualitative difference between graphs. While pooling based on spectral properties has been investigated extensively [9, 25], alternative geometric invariants like curvature or persistent homology have only been explored more recently [44, 20] and have not yet been applied to edge pooling specifically.
2.2 The Magnitude and Spread of Metric Spaces
Magnitude is an invariant of (finite) metric spaces that measures the ‘effective size’ of a space. It is a measure of entropy and diversity [30] that has first been proposed in theoretical ecology [39]. Since its mathematical formalisation by Leinster [28], magnitude has been connected to numerous key geometric invariants, such as entropy, curvature, density, volume, and intrinsic dimensionality [30]. Because of its intriguing theoretical properties magnitude has received increasing interest for machine learning tasks [31]. However, the magnitude of graphs [29], despite being a strong graph invariant, has not yet found its way into applications.
We consider an undirected finite graph $G=(X,E)$ with $n$ nodes as a finite metric space consisting of the node set $X$ equipped with a metric $d\colon X\times X\rightarrow\mathbb{R}_{\geq 0}$. Its similarity matrix $\zeta_{X}\subseteq\mathbb{R}^{n\times n}$ is defined by $\zeta_{X}(x,y)=e^{-d(x,y)}$ for $x,y\in X$. This allows us to introduce similarity-dependent notions of the diversity of metric spaces. To this end, we define a weighting on the metric space $(X,d)$, which is a vector $w\in\mathbb{R}^{n}$ such that $\zeta_{X}w=\mathds{1}$, where $\mathds{1}$ is the column vector of ones. Whenever such a weighting exists, the magnitude of the metric space $(X,d)$ is uniquely defined by $\text{Mag}(X)=\sum_{i=1}^{n}w(i)$. This is guaranteed if $\zeta_{X}$ is positive definite, which essentially means that $\zeta_{X}$ is invertible. A finite metric space with positive definite similarity matrix is called positive definite [33]. Metric spaces of negative type are positive definite [30]; this includes $\mathbb{R}^{n}$ equipped with the Euclidean distance [28], effective resistance distances[14], and diffusion distances [12].
Throughout this paper, we will refer to the magnitude of a graph $G$ as the magnitude of its associated metric space $(X,d)$. That is, we define the magnitude of a graph as
[…]
Closely related to magnitude, the spread of a metric space is another measure of ‘size’ introduced by Willerton [41]. Given a metric graph $G=(X,E)$ with the graph metric $d$, its spread is defined by
$\text{Sp}(G):=\sum_{x\in X}\frac{1}{\sum_{y\in X}e^{-d(x,y)}}.$ (2)
As diversity measures on graphs, both magnitude and spread summarise the number of distinct sub-communities in a network based on the distance metric and degree of similarity between nodes. This view on structural diversity naturally aligns with our goal of contracting redundant graph structures during pooling. Throughout this work, we investigate to what extent spread is a valid alternative to magnitude in practice. Magnitude is well-studied and has received increasing interest in both theory [30] and application [31]. However, computing it either requires inverting a matrix, solving a system of linear equations [31], or approximating [5], which can be notoriously expensive or unstable to compute. Metric-space spread in comparison can be computed given any distance (obviating the requirement of metric spaces of negative type), making it much more versatile [41]. Moreover, as the sum of reciprocal mean similarities, spread can be calculated or approximated [18] much more efficiently than magnitude and does not require inverting a matrix. Although spread has been studied less extensively [41, 18], there are strong reasons to assume that it shares the same advantages as magnitude. In fact, for a positive definite metric space $X$, we have $\mathrm{Sp}(X)\leq\mathrm{Mag}(X)$ [41, Theorem 2]. Moreover, magnitude and spread coincide for finite homogeneous metric spaces [41, Theorem 3], such as the ring graph in Figure 1. In practice, as we further explore in Section D.1, the magnitude and spread of graphs from real-world datasets, such as NCI1, exhibit nigh-perfect correlation when computed based on diffusion distances, underlining the strong connection between the two quantities.
2.3 Diffusion Geometry on Graphs
In this work, we use diffusion distances to compute magnitude and spread on graphs. This is motivated by their desirable theoretical properties and the capability of diffusion to aid and act along message passing in GNNs [23]. Diffusion operators are closely associated to random walks and are very efficient at identifying important structures in complex geometries while preserving local and non-linear structures [23]. The key idea is that the eigenvectors of the Markov matrices can be thought of as coordinates for the underlying graph structure [12]. This provides a vector space representation of the graph that can assess the dissimilarity between nodes.
We now briefly detail the type of diffusion distance used throughout this paper [12]. Consider a graph $G=(X,E)$ and its adjacency matrix $A$. The latter could be an affinity matrix derived from a kernel. Let $D$ be the diagonal degree matrix whose diagonal entries $D_{ii}=\sum_{j=0}^{n}A_{ij}$ equal the degree of each vertex. The symmetrically normalised graph Laplacian is defined as $\hat{L}=D^{-\frac{1}{2}}(D-A)D^{-\frac{1}{2}}$. This can be regarded as the simplest example of a Markov transition matrix and represents the probability of moving from one vertex to another. Since $\hat{L}$ is symmetric and positive definite it has positive eigenvalues $1=\lambda_{0}>\lambda_{1}>\lambda_{2}>\cdots\lambda_{N-1}\geq 0$ with eigenvectors $\{\psi_{l}\}_{l}$. This provides a natural embedding of the graph $G$ in Euclidean space given by:
$\Phi(x)=(\lambda_{1}\psi_{1}(x),\cdots,\lambda_{N-1}\psi_{N-1}(x))\text{ for }x\in X.$ (3)
The diffusion distance is defined by :
$d(x,y)=\left\lVert\Phi(x)-\Phi(y)\right\rVert_{2}\text{ for }x,y\in X.$ (4)
Theorem 1.
Any finite metric space $(X,d)$ endowed with the diffusion distance is positive definite.
Proof. See Section B.1.
As a consequence of Theorem 1, the magnitude of any metric graph equipped with this diffusion distance is well defined. Leveraging diffusion distances for our methods has further benefits. The normalised graph Laplacian, which works with the relative connectivity between nodes, is robust to varying node degrees and graph sizes. This ensures that diffusion distances are on comparable scales across graphs, which enables us to compute and compare magnitude and spread directly.
3 Methods
3.1 Magnitude-Guided Graph Pooling
At the heart of our approach, we use magnitude or spread to monitor and control structural changes in the graph during edge contraction pooling. Edge contraction is chosen as a pooling operation because it respects graph connectivity, while outputting a sparsely-connected graph representation. Conceptually, edge pooling thus aligns well with the goal of making minimal changes to the graph and keeping its diversity and geometry as unchanged as possible. Intuitively, pooled graphs with comparable magnitude will be similar in terms of effective size. That is, the effective number of distinct communities in the graphs are deemed similar based on their diffusion distance i.e. the information-flow between vertices.
Figure 2: Illustrating our proposed pooling method, MagEdgePool, on a graph from the ENZYME dataset across varying pooling ratios. Each edge is coloured by its magnitude difference, which measures the impact its contraction would have on the graph’s structural diversity. Edges with low magnitude differences are most redundant for the graph’s geometry and are collapsed first.
More formally, let $G=(V,E)$ be a graph and denote by $G/{e}$ the graph resulting from the contraction of the edge $e\in E$ in $G$. We choose a pooling ratio $r\in(0,1]$ and aim to reduce the graph to the corresponding number of nodes i.e. $k=\lfloor r\cdot|V|\rceil$. Initially, we set the pooled graph $G^{\prime}=(V^{\prime},E^{\prime}):=G$. To assess which edges to contract first, we determine their importance for the graph’s global geometry by computing a selection score for each edge defined as
$s(e)=|\text{Mag}(G)-\text{Mag}(G/{e})|.$ (5)
That is, we calculate the difference in magnitude between the original graph and the graph for which the edge has been collapsed. In this manner, we score an edge’s relevance for the graph structure by the impact its collapse would have on the graph’s magnitude. At each iteration $i$, we then select the edge with the lowest magnitude difference
$e_{i}\in\arg\min_{e\in E^{\prime}\setminus E_{c}}s(e)$ (6)
and assign $G^{\prime}=G^{\prime}/e_{i}$ where $E_{c}\subseteq E^{\prime}$ is the set of all edges that are not adjacent to an already contracted edge. The edge to contract $e_{i}$ is chosen randomly amongst the possible choices whenever there is more than one valid option. If all edges that meet this requirement have been collapsed, but the pooling ratio is not reached, we re-compute the edge scores on the new graph and repeat the same procedure. We stop if the pooling ratio is reached, i.e. $|V^{\prime}|=\lfloor r\cdot|V|\rceil$, or if there are no edges left in the reduced graph, i.e. $E^{\prime}=\emptyset$. This approach allows us to flexibly reduce graphs to any desired size.
Edge pooling then gives a hard assignment of nodes, where each vertex is assigned to a single super-node in the output graph based on the neighbours it has been merged with. To compress the node features, we average the features of any node that contributed to a pooled super-node. This ensures that information on all nodes’ features is preserved during pooling. Restricting the number of times a vertex can be merged is enforced to prevent our methods from collapsing entire portions of a graph early in the pooling process. This enables a more uniform pooling across the graph, which aids feature preservation and expressivity.
Our methods assume that the most redundant edges will be those whose removal would change the diversity of the graph the least. Magnitude informs us about the global importance of an edge. Intuitively, that means we want to start with merging edges from well connected communities whenever their contraction does not notably change the graph’s effective size. As illustrated in Figure 2, this ensures that globally influential edges will be collapsed late in the pooling progress. For example, the edge which bridges the two parts of this enzyme graph, is scored more highly and thus merged later than well-connected cliques. Redundant local structures are collapsed first, and the overall diffusion geometry is respected. For the graph in Figure 2 this ensures that key topological characteristics, such as the cycle, are retained across the pooling process. Our algorithm is explained in more detail in Section C.4. Whenever we use magnitude to compute the edge scores, we denote our algorithm by MagEdgePool. Similarly, the use of spread is denoted by SpreadEdgePool.
3.2 Theoretical Analysis
We now present theoretical properties of our pooling methods. First, we highlight fundamental invariances and properties of magnitude and spread used to design our algorithms. Further, we provide a bound on the difference in magnitude during pooling by the difference in spread, demonstrating the close relationship between the two. For a complete list or theorems and proofs see Section B.3.
Additivity for disjoint graphs.
An important property of magnitude is that it behaves like the cardinality of sets. This behaviour is useful when computing the magnitude of a disconnected graph by splitting the problem into calculating the magnitude of the disconnected components.
Theorem 2.
Consider a positive definite metric graph $G=G_{1}\sqcup G_{2}$ consisting of the disjoint union of two graphs $G_{1}$ and $G_{2}$. Then $\text{Mag}(G)=\text{Mag}(G_{1})+\text{Mag}(G_{2})$.
Isomorphism invariance.
Our pooling layers are invariant under isometries of the input graph provided that the choice of edges to contract at each iteration is deterministic in the case that the edge scores coincide. This is a consequence of the following result.
Theorem 3.
If the metric graphs $G_{1}$ and $G_{2}$ are isometric then $\text{Mag}(G_{1})=\text{Mag}(G_{2})$ and $\text{Sp}(G_{1})=\text{Sp}(G_{2})$.
Edge contraction on graphs.
Edge contraction is the main operation of our pooling methods and can be considered as a map $f$ between graphs. The following result provides a sufficient condition to ensure that the morphism $f$ is compatible with the metric structure.
Theorem 4.
Consider an edge-contraction map $f:(G_{1},d_{1})\rightarrow(G_{2},d_{2})$ between positive definite metric graphs. If the map is Lipschitz (i.e. $d_{2}(f(v_{1}),f(v_{2}))\leq d_{1}(v_{1},v_{2})\ \forall v_{1},v_{2}\in X_{1}$) then $\text{Mag}(G_{2})\leq\text{Mag}(G_{1})$.
Starting from an initial metric graph $G=(X,E)$, a Lipschitz edge-contraction map $f$ yields a sequence of graphs $\{G_{i}\}_{i=1}^{k}$ where $k$ is the number of edges that have been contracted. This sequence can be constructed as described by our algorithms in Section 3.1. Note that it is not guaranteed that MagEdgePool and SpreadEdgePool yield the same sequence of hierarchically pooled graphs. For this reason, we will refer to the graphs resulting from the $k^{th}$ edge-contraction with magnitude and spread as $G^{(k)}$ and $\widetilde{G}^{(k)}$, respectively. Let $\Delta^{(k)}\text{Mag}(G)=|\text{Mag}(G^{(k-1)})-\text{Mag}(G^{(k)})|$ and let $\Delta^{(k)}\text{Sp}(G)=|\text{Sp}(G^{(k-1)})-\text{Sp}(G^{(k)})|$.
Bounding magnitude by spread.
We track the difference in magnitude and spread throughout the edge contraction process detailed above. This allows us to propose an inequality that describes the relation between the difference of magnitude and spread during pooling. The bound then demonstrate the close conceptual relationship between SpreadEdgePool and MagEdgePool.
Theorem 5.
Consider a positive definite metric graph $G$ with positive weights. Assume that the edge-contraction maps describing MagEdgePool and SpreadEdgePool induce distance decreasing surjections on the vertex sets. If $|\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k)})|\leq C\Delta^{(k)}\text{Sp}(G)$, then
$\Delta^{(k)}\text{Mag}(G)\leq 3C\Delta^{(k)}\text{Sp}(G).$
Computational Complexity.
The time complexity of our pooling methods is independent of the GNN and is dominated by the cost of computing the edge scores in Equation 5. Given a graph $G=(X,E)$, magnitude has time complexity $O(|X|^{3})$. In comparison, spread has time complexity $O(|X|^{2})$ and can be more efficiently approximated [18]. Spread thus offers a considerably faster alternative. Now, let $O(C_{d})$ be the time complexity of computing the metric on $G$. The time complexity of SpreadEdgePool is dominated by $O(|E|(C_{d}+|X|^{2}+\text{log}|E|))$ and MagEdgePool by $O(|E|(C_{d}+|X|^{3}+\text{log}|E|))$. We note that on large graphs, it is possible to speed up the distance computations further to ensure scalability. See Section B.2 for a full description of computational costs and Section D.2 for an empirical evaluation, which shows that our algorithm performs on-par with existing pooling methods across the datasets evaluated throughout this work.
4 Experimental Results
Across our experiments, we address the following three key tasks: (i) graph classification performance, (ii) graph structure preservation during pooling, (iii) and performance across varying pooling ratios.
4.1 Graph Classification
Table 1: Mean and standard deviation of the graph classification accuracy. The best performing model is marked in bold. All models that did not perform significantly different from the best model are coloured green. The rightmost column shows the mean rank of each pooling method across datasets.
Method ENZYMES PROTEINS Mutagenicity DHFR IMDB-B IMDB-M NCI1 NCI109 Mean Rank
No Pooling 87.3 ± 2.5 73.8 ± 0.8 80.1 ± 1.3 71.4 ± 1.9 69.7 ± 0.7 46.0 ± 0.7 76.5 ± 1.8 74.3 ± 2.0 -
MagEdge 91.5 ± 3.2 76.4 ± 3.9 77.5 ± 2.7 88.0 ± 3.8 72.4 ± 1.7 47.4 ± 1.7 72.7 ± 2.4 73.0 ± 3.3 2.4
SpreadEdge 92.8 ± 1.6 75.1 ± 3.1 76.0 ± 4.0 90.7 ± 3.8 71.8 ± 1.5 47.3 ± 1.7 73.4 ± 2.5 71.8 ± 1.8 3.0
NDP 92.2 ± 1.6 73.7 ± 3.9 73.4 ± 3.1 79.6 ± 4.4 73.3 ± 2.0 47.3 ± 2.5 70.6 ± 2.2 70.0 ± 2.2 3.6
Graclus 91.3 ± 3.7 76.6 ± 3.7 72.5 ± 2.0 64.4 ± 5.8 71.9 ± 1.5 49.3 ± 2.4 68.8 ± 1.4 69.5 ± 2.1 5.6
NMF 78.6 ± 8.0 73.0 ± 8.1 71.0 ± 4.7 66.5 ± 7.7 69.4 ± 2.5 43.3 ± 1.7 71.0 ± 3.7 72.0 ± 5.3 6.0
TopK 82.2 ± 7.5 73.2 ± 1.4 75.8 ± 4.7 68.9 ± 3.0 68.9 ± 1.5 45.6 ± 1.0 75.3 ± 2.4 73.9 ± 3.3 4.9
SAGPool 82.4 ± 4.5 73.8 ± 1.3 76.0 ± 2.6 69.9 ± 3.0 69.1 ± 0.6 45.7 ± 0.5 74.3 ± 2.8 74.0 ± 2.3 4.7
DiffPool 74.0 ± 5.7 68.9 ± 2.0 68.4 ± 1.9 79.8 ± 3.2 68.3 ± 0.8 44.4 ± 0.8 68.9 ± 1.0 68.3 ± 1.9 7.6
MinCut 80.2 ± 6.6 75.6 ± 1.3 70.9 ± 1.5 63.8 ± 3.7 69.3 ± 0.7 46.1 ± 0.8 66.7 ± 1.4 66.9 ± 2.0 7.4
A key aim of implementing graph pooling layers is to preserve task-relevant information while reducing computational costs. In particular, useful pooling layers guarantee good performance across a wide range of different datasets, thus capturing essential information for the task at hand. We thus investigate how well our aim to preserve structural diversity during edge pooling translates to practical performance at graph classification tasks.
Experimental Setup.
We evaluate 8 different graph datasets [35], as detailed in Section C.2. Whenever node features are not available, we use node degree as an input feature. To ensure a fair comparison with alternative pooling methods, we follow the experimental setup by Grattarola et al. [25] and guidance by Errica et al. [19] for model selection. Specifically, we plug in each pooling layer into the model architecture specified by Grattarola et al. [25], which is of the following form:
$\text{MLP}(\mathbf{X})\rightarrow\text{GNN}(\mathbf{X},\mathbf{A})\rightarrow\text{POOL}(\mathbf{X},\mathbf{A})\rightarrow\text{GNN}(\mathbf{X},\mathbf{A})\rightarrow\text{GlobalSum}(\mathbf{X})\rightarrow\text{MLP}(\mathbf{X})$
The model includes pre-processing and post-processing MLPs with 2 layers, 256 hidden units, ReLU activation, and batch normalization. $\text{GNN}(\mathbf{X},\mathbf{A})$ refers to a graph neural network layer, more specifically a general convolutional layer [46] with parameters chosen according to the best results achieved by You et al. [46]. As an intermediate layer, $\text{POOL}(\mathbf{X},\mathbf{A})$ corresponds to a specific pooling layer. All pooling layers are configured to pool each graph to around 50% of nodes. We also compare with ‘No Pooling,’ the same model architecture without any pooling layers. We use 10-fold stratified cross-validation and further partition the training data into 90% training and 10% validation data while keeping the labels balanced between splits. Finally, we report the best test accuracy of each model trained using Adam with a cross-entropy loss (batch size 32, learning rate $0.0005$, and early stopping based on the validation loss with a patience of 50 epochs). Further details are described in Section C.5.2.
Classification Results.
Table 1 reports the mean and standard deviation of the test accuracy achieved by different pooling methods. We furthermore highlight which methods do not perform statistically significantly different from the best model (using pairwise Wilcoxon signed-rank tests applied to the accuracy scores and employing Holm–Bonferroni correction at a significance threshold of $p=0.05$), thus permitting us to identify pooling methods that achieve top performance. Notably, both MagEdgePool and SpreadEdgePool achieve the best mean ranks across datasets in terms of their accuracy. Further, they are always among the top-performing methods across all evaluated datasets. Altogether, both their ranking and their individual accuracy scores thus demonstrate superior and consistently high performance across graph classification tasks. The performance benefits of our proposed pooling methods are most pronounced on DHFR [40], where our methods surpass even the GNN without pooling layer by around 17 percentage points. This provides evidence that the regularising effects of our pooling approach can help reduce overfitting, especially for small datasets and geometrically-rich graphs. For other biological datasets (Mutagenicity, NCI1 and NC109), our methods show competitive performance with trainable layers, indicating that the introduction of additional trainable components into the pooling layer is not necessary to guarantee high task performance. On ENZYMES, DHFR, IMDB-BINARY, and IMDB-MULTI, non-trainable methods generally outperform trainable pooling layers, with MagEdgePool and SpreadEdgePool consistently reaching high accuracy.
Comparing pooling methods to using no pooling, we observe that MagEdgePool and SpreadEdgePool reach similar or even higher performance across datasets. For six datasets, our diversity-guided pooling methods improve mean accuracy, indicating that pooling retains task-relevant information while aiding the generalisation capabilities of the GNN. MagEdgePool and SpreadEdgePool act as interpretable and expressive pooling transformations, which reduce the computational costs making GNNs learn from graphs’ coarsened geometry. As reported in Table 1, MagEdgePool and SpreadEdgePool achieve very similar accuracies across datasets superior to alternative pooling layers. In practice, especially for large graphs, we recommend using SpreadEdgePool due to its high predictive performance and superior computational efficiency.
4.2 Magnitude and Graph Structure Preservation
Figure 3: Structure preservation for all graphs in the NCI1 dataset across varying pooling ratios. Left: The spectral distance between the normalised Laplacians of the the original and the pooled graphs. Right: The relative difference in magnitude, which summarises the proportional difference in structural diversity after pooling. Violinplots show the variability across graphs at pooling ratio 0.5.
Motivated by the visual comparison of graphs pooled using different pooling layers from Figure 1, we next set out to investigate the link between structure preservation and task performance. Specifically, we choose NCI1, a dataset of 4,110 graphs corresponding to chemical compounds, because it has been shown to possess both informative features and task-relevant graph structures [13]. We follow the same classification procedure as before and extract the pooled graph representations after training the GNNs described in Table 1. Three pooling layers, MinCut, DiffPool and NMF, were removed from further comparison because they showed notably worse qualitative results for the motivating examples in Figure 1 and classification performance in Table 1. NDP and Graclus are evaluated across fewer pooling ratios than more adaptive methods, because they pool graphs to around half their size at every step. To assess graph structure preservation we use the spectral distance defined as $\sqrt{\sum_{k=1}^{K}(\lambda_{k}-\lambda_{k}^{\prime})^{2}}$, the $l_{2}$-norm between the eigenspectra of the normalised Laplacians of the original and the pooled graphs [42]. We also report the magnitude difference between graphs to evaluate the preservation of structural diversity.
MagEdgePool and SpreadEdgePool show low spectral distances across pooling ratios as visualised in Figure 3. This indicates that contracting edges guided by structural diversity preserves key spectral properties. Our methods also demonstrate low magnitude differences confirming that they perform as intended. In fact, the structure preservation scores for MagEdgePool and SpreadEdgePool coincide almost perfectly, giving empirical evidence that spread offers an alterative to magnitude. Figure 3 indicates that preserving magnitude corresponds to lower spectral distances and better retention of spectral properties. This link supports our motivation of guiding pooling by magnitude.
Alternative pooling layers fail to effectively preserve graphs’ structural properties during both qualitative and quantitative comparisons to varying extents. Node decimal pooling (NDP) [9] was specifically designed to preserve spectral properties during pooling. However, it still reaches both higher spectral distances and higher magnitude differences than MagEdgePool on average across pooling ratios, as visualised in Figure 3. Finally, the sparse pooling layers Graclus, TopKPool, and SAGPool, all show higher spectral distances than our approach. This difference is even more pronounced in terms of magnitude differences, where all these three methods demonstrate high distortion of the underlying metric space diversity. These findings agree with the qualitative comparisons between graphs pooled using different pooling layers as visualised in Figure 1. We thus conclude that MagEdgePool and SpreadEdgePool successfully encode graphs’ coarsened geometry during pooling, surpassing alternative pooling methods.
4.3 Pooling Ratio and Task Performance
Figure 4: Classification performance across varying the pooling ratio for different pooling layers. Pooling is applied as part of a GIN architecture. Results are shown for the ENZYME and NCI1 datasets. Lines show the mean and shaded areas the standard deviation of the test accuracy.
From Table 1 we observe that it is possible to reach very high performance on benchmark datasets even while pooling each graph to half its size. Based on this, we further investigate how the pooling ratio influences pooling layers and their classification performance. We consider two datasets, NCI1 and ENZYMES, which contains 600 graphs representing protein tertiary structures from 6 classes of enzymes and is selected as an example of a multi-class prediction task. We keep the same experimental setup described in Table 1, but use GIN layers instead of general convolutional layers to further assess whether the trends in performance differ across model architectures. Figure 4 then reports the classification accuracy achieved by each pooling layer for varying pooling ratios.
Notably, we observe that MagEdgePool and SpreadEdgePool consistently reach very high test accuracies even at low pooling rations. Meanwhile, the performance of other non-trainable methods drops notably more when graphs are pooled to up to 6.25% of their original size showing that they fail to preserve task-relevant information for both ENZYMES and NCI1. We note that the trainable pooling layers, TopK and SAGPool, reach higher or comparable performance on NCI1 for pooling ratios above 50%, but decrease in accuracy for more extreme pooling ratios. They consistently perform worse on ENZYMES, indicating that they distort important graph features or key graph structures during pooling. MagEdgePool and SpreadEdgePool in comparison reach superior performance and lower decreases in accuracy across varying pooling ratios demonstrating their potential to offer reliable, interpretable and stable pooling operations. Overall, the reported accuracies in Figure 4 agree with results in Table 1 indicating that our observations hold for varying choices of GNN layers. We thus find across experiments that MagEdgePool and SpreadEdgePool constitute useful general-purpose pooling approaches, demonstrating their capability to faithfully encoding graphs’ geometry, which ensures stable performance across pooling ratios, datasets and GNN architectures.
5 Discussion
Limitations.
We implicitly require redundancy and homophily in the graph representation which can be exploited via geometry-guided pooling. Our approaches further assume that preserving graph structure is beneficial to the learning task. Otherwise, edge pooling still aggregates features faithfully, but the geometric objective might not be necessary to ensure high performance. Moreover, our algorithm relies on efficient distance computations and we only explore one specific case of diffusion distances, which could be generalised further in future work. Although the non-trainable nature of our pooling methods can be limiting, our experiments nevertheless demonstrate that trainable pooling layers do not guarantee higher performance on standard benchmark datasets for graph classification.
Conclusions.
Across experiments we find that MagEdgePool and SpreadEdgePool constitute useful general-purpose pooling approaches. Our methods are competitive compared to state-of-the-art pooling layers for graph classification tasks and perform well for a wide range of datasets and pooling ratios. We find that guiding edge pooling to preserve graphs’ structural diversity successfully encodes key graph properties and ensures stable and interpretable performance. Further, we overcome one major limitation of computing magnitude on large graphs by proposing the spread of a metric space as a much faster and closely-related alternative, which has the potential to aid research in geometric deep learning and efficient diversity evaluation beyond the scope of this paper. For future work, we find that there is strong potential to apply our methods with alternative graph distances, or to determine an ideal pooling ratio automatically using magnitude or spread.
Acknowledgements
This work was partially funded by the Helmholtz International Lab Causal Cell Dynamics InterLabs-0029 grant by the Initiative and Networking Fund of the Hermann von Helmholtz-Association Deutscher Forschungszentren e.V. [K.L.], Mila EDI scholarships [L.M.], Humboldt Research Fellowship, CIFAR AI Chair, NSERC Discovery grant 03267, FRQNT grant 343567, and NSF grant DMS-2327211 [G.W.]. This work has received funding from the Swiss State Secretariat for Education, Research, and Innovation (SERI). K.L. is supported by the Helmholtz Association under the joint research school ‘Munich School for Data Science (MUDS).’ The content provided here is solely the responsibility of the authors and does not necessarily represent the views of the funding agencies.
Appendix A Technical Appendices and Supplementary Material
To elaborate on the results reported in our main paper, we first detail extended theoretical results and proofs for our theoretical contributions. In particular, we investigate computational complexity as well as the relationship between magnitude and spread. Next, we detail our experimental evaluation, the assets used for our experiments, and the algorithm describing our pooling methods. Finally, we report extended results on the experiments included in our main paper.
Appendix B Theoretical Analysis
This section details full proofs and extended explanations for the mathematical theory introduced in Section 2 and the theoretical analysis of our pooling methods described in Section 3.2.
B.1 Diffusion Distances
As detailed in Section 2.3, the diffusion distance is defined by:
$d(x,y)=\left\lVert\Phi(x)-\Phi(y)\right\rVert_{2}\text{ for }x,y\in X.$ (7)
See 1
Proof.
By definition of the diffusion distance, the map $\Phi:X\rightarrow\mathbb{R}^{N-1}$ in Equation 3 defines an isometry $(X,d)\hookrightarrow l_{2}^{N-1}:=(\mathbb{R}^{N-1},d_{2})$, where $d_{2}$ is the metric induced by the $l_{2}$-norm. Finally, by Theorem 2.5.3 in Leinster [28], subsets of Euclidean space $l_{2}^{N-1}$ are positive definite. ∎
B.2 Computational Complexity
We next analyse the computational complexity of our pooling methods, which are described in Section C.4 and Section 3.1. Specifically, we expand on the statements in Section 3.2 by detailing the computational complexity of the pooling process. Given a graph $G=(X,E)$, let $k=\lfloor(1-r)|X|\rfloor$ be the number of nodes that should be contracted as determined by the pooling ratio $r$. The time complexity of our pooling approach can be split up into the following steps:
Computing magnitude or spread.
Magnitude has time complexity $O(|X|^{3})$ and can further be approximated via iterative normalisation in $O(i\times|S_{i}|\times|X|^{2})$ time assuming $G$ has a positive weighting where $i$ is the number of iterations and $S_{i}\subset X$ [5]. Spread computations have time complexity $O(|X|^{2})$, which is a notable improvement to magnitude. It is possible to approximate spread computations via subsets [18] or iterative optimisation using mini-batching [5]. For $i$ iterations on subsets $S_{i}\subset X$, the time complexity of approximating spread reduces to $O(i\times|S_{i}|\times|X|)$ [18, 5]. Spread thus offers a much faster alternative to magnitude and can scale to large graphs considerably more efficiently.
Computing distances and similarities.
For large datasets, it is key to speed up the distance calculations. Diffusion distances have time complexity $O(|X|^{3})$, but can be reduced to $O(k|X|^{2})$ when restricting the computations to the top k eigenvectors. Diffusion maps can further be approximated via low-rank approximations. To reduce the cost of repeated distance computations, it is possible to approximate the metric on the reduced graph $G/e$ by directly updating the distances for $G$. Given a distance matrix, computing the similarity matrix then has linear time complexity in the number of entries. Denote the time complexity of computing the distances and similarities by $O(C_{d})$.
Edge contraction.
To get $G^{\prime}=G/e$, contracting an edge $e\in E$ takes $O(|X|)$ time.
Edge score computations.
For each edge, its edge score is computed by applying the edge contraction and computing the magnitude or spread of the reduced graph, which takes $O(|X|+C_{d}+C_{S})$ time, where $C_{S}$ refers to the cost of computing either magnitude or spread as detailed above. Note that the computation of these edge scores is independent across edges and can be parallelised.
Edge score sorting.
The edge scores can be sorted from lowest to highest in $O(|E|\log|E|)$ time.
Feature aggregation.
The node features, $\mathbf{F}\subseteq\mathbb{R}^{|X|,F}$, can be aggregated in $O(|X|\times F)$ time.
Putting this all together, we get that the cost of our pooling algorithms can be described by a worst-case time complexity of
$O(|E|(|X|+C_{d}+C_{S}+log|E|)+|X|(F+k))$
if $k\leq 0.5|X|$ and the graph is not pooled to less than half its size. Otherwise, if $k>0.5|X|$, we re-compute the edge scores whenever no valid edges are left as described in Section 3.1. In this scenario, the first term of the complexity expression is repeated, corresponding to re-computations on successively smaller graphs. In summary, the overall time complexity of our pooling method is dominated by the cost of calculating and sorting the edge scores. This cost is independent of the choice of GNN architecture, ensuring that the training costs remain stable and do not escalate with model complexity. In practice, as further explored in Section D.2, we thus find that our pooling algorithms perform on par with alternative pooling layers in terms of computational efficiency.
B.3 Magnitude and Spread
B.3.1 Additivity for disjoint graphs
As a measure of the effective size, one appealing property of magnitude is that it behaves akin to cardinality. In fact, magnitude is additive when taking the disjoint union of multiple metric spaces.
See 2
Proof.
Let $G=(X,E)$ together with the metric $d$ be a positive definite metric graph. Assume $G$ is a disjoint union of two metric graphs $(G_{1},d_{1})$ and $(G_{2},d_{2})$ over the vertex sets $X_{1}$ and $X_{2}$, respectively, such that $d_{|_{X_{1}}}=d_{1}$ and $d_{|_{X_{2}}}=d_{2}$ and $d(x_{1},x_{2})=\infty$ for all $x_{1}\in X_{1},\ x_{2}\in X_{2}$. Then, $\zeta_{G}=\zeta_{G_{1}}\oplus\zeta_{G_{2}}$ and $\text{Mag}(G)=\text{Mag}(G_{1})+\text{Mag}(G_{2})$. ∎
This result applies to graphs equipped with the shortest-path distance or the diffusion distance considered in this paper, because the distance between two nodes depends only on the connected component they belong to and is infinite if there is no path between them. Therefore, we can naturally see the similarity matrix $\zeta_{G}$ as block-diagonal and compute the magnitude of $G$ by summing up the magnitude of its disconnected subgraphs.
B.3.2 Isomorphism invariance
A key property of magnitude and spread is that they are isometry invariants of metric spaces. Note that by graph isometry we mean an isometry on the underlying vertex set equipped with a metric.
See 3
Proof.
Let $f:(X_{1},d_{1})\rightarrow(X_{2},d_{2})$ be a bijective isometry between the metric graphs $G_{1}$ and $G_{2}$, respectively. Then, for all $x,y\in X_{1}$ we have that $d_{2}(f(x),f(y))=d_{1}(x,y)$. A consequence of the bijectivity of $f$ is that the distance matrices coincide (up to permutations) and that $\zeta_{1}=\zeta_{2}$. This implies in turn that $\text{Mag}(G_{1})=\text{Mag}(G_{2})$ and that $\text{Sp}(G_{1})=\text{Sp}(G_{2})$. ∎
Based on this property for magnitude and spread, we can show that isometry invariance also holds for our proposed pooling algorithm further detailed in Section C.4.
Corollary 1.
MagEdgePool and SpreadEdgePool are isometry-invariant if applied to isomorphic graphs provided the choice of edges to contract at each iteration is deterministic whenever the edge scores coincide.
Proof.
Let $f:(G_{1},d_{1})\rightarrow(G_{2},d_{2})$ be a graph isomorphism and an isometry. From Theorem 3 we have that the edge scores as defined in Equation 5 for $e_{1}\in E_{1}$ and $f(e_{1})\in E_{2}$ will coincide, i.e. $s(e_{1})=s(f(e_{1}))$. Because the choice of edges to contract at each iteration is further assumed to be deterministic if edge scores coincide, it follows that every edge to contract in $G_{2}$ corresponds to the image $f(e)$ of an edge $e$ to contract in $G_{1}$ and vice versa. Hence, the pooled graphs output by our algorithm are isomorphic. ∎
B.3.3 Edge contraction on graphs
We will first recall important results about magnitude in the context of (strictly) positive definite finite metric spaces and refer the interested reader to Leinster [28] for further details.
Proposition 1 (Proposition 2.4.3, [28]).
Let $(X,d)$ be a positive definite metric space with finite cardinality $|X|=n$. Then
$\text{Mag}(X)=\sup_{v\in\mathbb{R}^{n}\backslash\{0\}}\frac{(\sum_{i=0}^{n}v_{i})^{2}}{v^{t}\zeta_{X}v}.$ (8)
Proposition 1 implies that the magnitude of positive definite metric spaces is always positive. Another consequence of this result is a monotonicity property on subsets of these metric spaces.
Corollary 2 (Corollary 2.4.4, [28]).
Let $(X,d)$ be a positive definite finite metric space and consider a subset $Y\subset X$ (endowed with the induced metric). Then
$\text{Mag}(Y)\leq\text{Mag}(X).$ (9)
We will now show an analogous result for graphs constructed via edge contraction.
See 4
Proof.
Let $f:(G_{1},d_{1})\rightarrow(G_{2},d_{2})$ be an edge-contraction map and let $n:=|X_{1}|$ and $m:=|X_{2}|$. We will identify $\mathbb{R}^{m}$ with a subset of $\mathbb{R}^{n}=\mathbb{R}^{m}\oplus\mathbb{R}^{n-m}$ using the map $(v_{1},\cdots,v_{m})\in\mathbb{R}^{m}\hookrightarrow(v_{1},\cdots,v_{m},0,\cdots,0)\in\mathbb{R}^{n}$. Then,
$\begin{split}v^{t}\zeta_{X_{1}}v&=\sum_{i,j}v_{i}\zeta_{X_{1}}[i,j]v_{j}\\ &=\sum_{i,j}v_{i}\big{(}e^{-d_{1}(x_{i},x_{j})}\big{)}v_{j}\\ &\leq\sum_{i,j}v_{i}\zeta_{X_{2}}[i,j]v_{j},\end{split}$
and
$\frac{1}{v^{t}\zeta_{X_{2}}v}\leq\frac{1}{v^{t}\zeta_{X_{1}}v}\ \forall\ v\neq 0.$ (10)
Finally, by Proposition 1 and Inequality 10, we get that
$\begin{split}\text{Mag}(G_{1})&=\sup_{v\in\mathbb{R}^{n}\backslash\{0\}}\frac{(\sum_{i=0}^{n}v_{i})^{2}}{v^{t}\zeta_{X_{1}}v}\geq\sup_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{(\sum_{i=0}^{n}v_{i})^{2}}{v^{t}\zeta_{X_{1}}v}\\ &\geq\sup_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{(\sum_{i=0}^{m}v_{i})^{2}}{v^{t}\zeta_{X_{2}}v}=\text{Mag}(G_{2})\end{split}$
B.3.4 Bounding magnitude by spread
Recall that for positive definite metric spaces, magnitude is known to be an upper bound for spread.
Theorem 6 (Theorem 2.2. [41]).
Suppose that $X$ is a finite metric space. If $X$ is positive definite then
$\text{Sp}(X)\leq\text{Mag}(X).$
We will now use this bound as well as the results in Section B.3.3 to investigate the relationship between MagEdgePool and SpreadEdgePool. Through a process of iterated edge contraction, our pooling algorithm produces a sequence of hierarchically pooled graphs (as described in Section 3.1 and Section C.4). Note that it is not guaranteed that MagEdgePool and SpreadEdgePool yield the same sequence. For this reason, we will refer to the graphs resulting from the $k^{th}$ edge-contraction with magnitude and spread by $G^{(k)}$ and $\widetilde{G}^{(k)}$ respectively.
For each $k$ the edge contraction map $G^{(k)}\rightarrow G^{(k+1)}$ is a surjection on the underlying vertex sets $X^{(k)}$ and $X^{(k+1)}$ respectively, i.e $X^{(k+1)}\subset X^{(k)}$. Moreover, we will assume that for any $k$ this map is distance-decreasing. That is, $d^{(k+1)}(f(x_{i}),f(x_{j}))\leq d^{(k)}(x_{i},x_{j})$ for all $x_{i},x_{j}\in X^{(k+1)}$.
Recall that for any $k$, $\text{Mag}(G^{(k)})$ is the magnitude of a finite positive definite metric space $(X^{(k)},d^{(k)})$. Then, by Theorem 4, we deduce that for any $k$,
$\text{Mag}(G^{(k+1)})\leq\text{Mag}(G^{(k)})$ (11)
Note that by construction, scoring the edges in Algorithm C.4 translates into the following:
$X^{(k)}=\text{argmin}_{Y\subset X^{(k-1)},|Y|+1=|X^{(k-1)}|}|\text{Mag}(X^{(k-1)})-\text{Mag}(Y)|$ (12)
and,
$\widetilde{X}^{(k)}=\text{argmin}_{Y\subset\widetilde{X}^{(k-1)},|Y|+1=|\widetilde{X}^{(k-1)}|}|\text{Sp}(\widetilde{X}^{(k-1)})-\text{Sp}(Y)|.$ (13)
Then, by monotonicity of magnitude (Inequality 11),
$X^{(k)}=\text{argmax}_{Y\subset X^{(k-1)},|Y|+1=|X^{(k-1)}|}\text{Mag}(Y)=\text{argmax}_{Y\subset X^{(k-1)}}\text{Mag}(Y).$
Let $\Delta^{(k)}\text{Mag}(G)=|\text{Mag}(G^{(k-1)})-\text{Mag}(G^{(k)})|$ and let $\Delta^{(k)}\text{Sp}(G)=|\text{Sp}(G^{(k-1)})-\text{Sp}(G^{(k)})|$. For the following result, we assume that scores are only computed once and that they are the only criterion for edge contraction. Furthermore, we will assume that spread is monotonically decreasing.
See 5
Proof.
For any $k$ we have the following inequality:
$\begin{split}|\text{Mag}(G^{(k-1)})-\text{Mag}(G^{(k)})|&\leq|\text{Sp}(G^{(k-1)})-\text{Sp}(G^{(k)})|+|\text{Mag}(G^{(k)})-\text{Sp}(G^{(k)})|\\ &+|\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k-1)})|.\end{split}$
Assume that $|\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k)})|\leq C\Delta^{(k)}\text{Sp}(G)$ for some constant $C>0$. By the monotonicity of magnitude (Theorem 4), we get that $\text{Mag}(G^{(k)})\leq\text{Mag}(G^{(k-1)})$ and,
$\text{Mag}(G^{(k)})-\text{Sp}(G^{(k)})\leq\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k)})\leq C\Delta^{(k)}\text{Sp}(G).$
Similarly, assuming monotonicity of spread yields $\text{Sp}(G^{(k)})\leq\text{Sp}(G^{(k-1)})$ and,
$\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k-1)})\leq\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k)})\leq C\Delta^{(k)}\text{Sp}(G).$
Since $0\leq\text{Sp}(G^{(k-1)})-\text{Sp}(G^{(k)})\leq\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k)})$, the constant $C$ must be greater than or equal to 1. We conclude that
$\begin{split}\Delta^{(k)}\text{Mag}(G)\leq 3C\Delta^{(k)}\text{Sp}(G).\end{split}$
Appendix C Extended Methods
C.1 Hardware and Software
The experiments reported in our study were implemented using spektral 1.3.1 [24]***https://graphneural.network/ available under an MIT license., and tensorflow 2.16.2 [1]https://pypi.org/project/tensorflow/2.16.2/ available under the Apache Software License (Apache 2.0).. As further detailed in Section C.5.2 and Table 1, we base our graph classification experiments on the benchmark setup and code by Grattarola et al. [25]https://github.com/danielegrattarola/SRC available to the research community (Grattarola et al. [25])., which also include implementations for the pooling layers compared across our study. By relying on this existing framework, we aim to ensure the reproducibility of our results.
Further, to calculate magnitude and spread we rely on magnipy, a Python package by Limbeck et al. [31]§§§https://github.com/aidos-lab/magnipy available under a BSD 3-Clause License. for magnitude and diversity computations. Further, as a novel contribution of this paper, we extend the computation of magnitude to graph data and novel graph metrics. Specifically, we modify the computations, so that the magnitude of disconnected subgraphs is computed separately (based on Theorem 2) using the NetworkXhttps://github.com/networkx/networkx available under a BSD 3-Clause License. package. We also implement graph distances that have not previously been used to compute magnitude, such as the diffusion distances detailed in Section 2.3. Further details on the code for implementing our proposed pooling methods and our experiments can be found in our supplementary code submission.
All experiments were conducted on a high performance cluster with hardware specifications as detailed in Table 2. In particular, all experiment were run requesting one GPU with 32GB video memory or less.
Table 2: Summary of the compute resources used for our experiments.
Inventory Models
Available CPUs Intel Xeon (Gold 6128, 6130, 6134, 6136, 6142, 6240, 6248R)
Intel Xeon Platinum (8280L, 8480+, 8468, 8562Y+)
Intel Xeon (E7-4850, E5620, 4114, 6126)
AMD EPYC (7262, 7413, 7513, 7713, 7742)
AMD Opteron (6128, 6164 HE, 6234, 6272, 6376 x2)
Available GPUs NVIDIA Tesla (K80, P100, V100)
NVIDIA A100 (20GB, 40GB, 80GB PCIe)
NVIDIA H100 (80GB PCIe)
NVIDIA Quadro RTX 8000
C.2 Datasets
We briefly describe the graph datasets analysed throughout our work. Simulated graphs, as used for Figure 1, are created using either PyGSP https://pygsp.readthedocs.io/en/stable/ available under a BSD-3-Clause license. or NetworkX******https://github.com/networkx/networkx available under a BSD 3-Clause License. and all example graphs were created to consist of 64 nodes.
For our main graph classification experiments, we analyse six graph datasets taken from biological or chemical applications [40, 10, 37, 17, 38], and two datasets which represent social networks [43]. All datasets are taken either from the TUDataset††††††https://chrsmrrs.github.io/datasets/ available under a CC BY 4.0 license. benchmark [35] or the Open Graph Benchmark‡‡‡‡‡‡https://ogb.stanford.edu/ available under an MIT licence..
More specifically, the main results in Table 1 analyse the following graph classification datasets described in Table 3. Note that we only consider node and not edge features for our experiments.
Table 3: Summary of the graph datasets considered for our experiments.
dataset library # classes # node features # graphs avg # nodes avg # edges
MUTAG TUDataset 2 no 187 18 40
Enzymes TUDataset 6 18 600 33 62
COX2 TUDataset 2 3 467 41 44
DHFR TUDataset 2 3 756 42 45
IMDB-B TUDataset 2 no 1000 20 97
IMDB-M TUDataset 3 no 1500 13 65
AIDS TUDataset 2 4 2000 15 16
Proteins TUDataset 2 29 1113 39 72
Mutagenicity TUDataset 2 no 4337 30 31
NCI1 TUDataset 2 no 4110 30 32
NCI109 TUDataset 2 no 4127 30 32
OGBG-MOLHIV OGB 2 9 41127 25 27
BZR TUDataset 2 3 405 36 38
BZR_MD TUDataset 2 no 306 21 225
COX2_MD TUDataset 2 no 303 26 335
DHFR_MD TUDataset 2 no 393 24 283
ER_MD TUDataset 2 no 446 21 235
C.3 Magnitude and Spread Computations
Across our experiments, we compute magnitude and spread as outlined in the main text, implemented in our code submission, and further described in Section C.1. Elaborating on these descriptions, we now aim to give an extended explanation of practical and theoretical considerations for computing the magnitude of graphs in practice.
Defining the magnitude of a graph.
In mathematical literature, the magnitude of graphs is often studied with the shortest path metric [29]. However, shortest path distances are not guaranteed to be of negative type, thus leading to scenarios and well-known examples for which the similarity matrix is not invertible and magnitude based on this metric cannot be computed [28]. In comparison, resistance distances, diffusion distances, or Euclidean distance always permit the computation of magnitude. Because of this difference in the choice of distance metric, we note that our definition of the magnitude of a graph in Section 2.2 differs from the definition used by Leinster [29].
Magnitude and spread as multi-scale functions.
Note that magnitude and spread can also be defined as multi-scale functions i.e. $t\mapsto\text{Mag}((X,t\cdot d))$ for a metric space $(X,d)$ and a scale parameter $t\in\mathbb{R}^{+}$. This parameter $t$ can be likened to choosing a kernel bandwidth or the scale of distances or similarity determining when observations are considered to be distinct. In practical applications, it is advisable to carefully consider the choice of scaling factor $t$ or the type of normalisation used to compare distances [31]. Limbeck et al. [31] propose a heuristic that uses root-finding to find a suitably large $t$. However, this requires repeated computations of magnitude and increases the computational costs. A faster and more desirable default choice of $t$ would be based solely on the distance metric. For distances that are not otherwise scaled or normalised, we therefore recommend the usage of faster heuristics, such as the median heuristic for choosing the kernel-bandwidth i.e. the scale parameter $t$ [22]. For diffusion distances, we find that setting $t=1$ is sufficient for our goals. This is because, as discussed in Section 2.3, diffusion distances are computed from the normalised graph Laplacians and are inherently comparable across graphs. Nevertheless, investigating magnitude and spread as multi-scale functions on graphs remains an interesting extension for further work.
Magnitude and spread as diversity measures.
We extensively discuss the relationship between magnitude and spread throughout our work. However, our main paper does not have the space to fully explain the theoretical motivations behind the formulations of magnitude, spread, and other generalised measures of diversity. For a more complete discussion we refer the interested reader to Leinster [30], an extensive reference work on the mathematical motivation behind entropy and diversity. Furthermore, Willerton [41] specifically discusses the spread of a metric space, and Limbeck et al. [31] describe the practical usage of magnitude as a diversity measure in ML. These works also give descriptions on how and why the magnitude or spread of a metric space can be interpreted as an effective size i.e. as the effective number of distinct points in a metric space or the number of dissimilar nodes in a graph.
C.4 Pooling Algorithm
We now detail our pooling algorithm introduced in Section 3.1 by describing a pseudocode implementation. Note that to describe the algorithm we assume we have pre-selected a distance metric for computing either magnitude or spread.
Algorithm 1 Graph Pooling Methods: SpreadEdgePool and MagEdgePool
0:  input graph $G=(X,E)$, node features $\mathbf{F}\in\mathbb{R}^{n\times d}$, pooling ratio $r\in(0,1]$, diversity measure $\text{Mag}(G)\text{ or }\text{Sp}(G)$
0:  Pooled graph $G^{\prime}=(X^{\prime},E^{\prime})$, pooled features $\mathbf{F}^{\prime}$
1:  Initialise the super-node set $\mathcal{S}(x)\leftarrow x$ for all $x\in X$
2:  Initialise the set of contracted edges $E_{c}\leftarrow\emptyset$
3:  Initialise the pooled graph $G^{\prime}\leftarrow G$
4:  Compute initial edge scores:
$s(e)=\left|\text{Mag}(G)-\text{Mag}(G/e)\right|\quad\forall e\in E$
5:  while $|X^{\prime}|\neq\lfloor r|X|\rceil$ AND $|E^{\prime}|\neq\emptyset$ do
6:     Select edge $e=(x,y)=\arg\min_{e\in E\setminus E_{c}}s(e)$
7:     if $e$ is not adjacent to any previously contracted edge in $E_{c}$ then
8:        Contract edge $e$, update $G^{\prime}\leftarrow G^{\prime}/e$
9:        Add $e$ to $E_{c}$
10:        Update the node selection: merge $\mathcal{S}(x)$ and $\mathcal{S}(y)$
11:     end if
12:     if no more valid edges AND pooling ratio not reached then
13:        Recompute the edge scores $s(e)$ on the updated graph $G^{\prime}$
14:        Reset $E_{c}\leftarrow\emptyset$
15:     end if
16:  end while
17:  Initialize $\mathbf{F}^{\prime}\leftarrow\emptyset$
18:  for each supernode representative $w\in\mathcal{S}$ do
19:     Let $S_{w}=\{x\in X\mid\mathcal{S}(x)=w\}$
20:     Compute aggregated feature:
$\mathbf{F}^{\prime}_{w,:}=\frac{1}{|S_{w}|}\sum_{x\in S_{w}}\mathbf{F}_{x,:}$
21:     Append $\mathbf{F}^{\prime}_{w}$ to $\mathbf{F}^{\prime}$
22:  end for
23:  return  Pooled graph $G^{\prime}$, pooled features $\mathbf{F}^{\prime}$
C.5 Extended Experimental Details
We briefly describe extended details on the experimental setup used for our main experiments.
C.5.1 Structure Preservation Experiment and Pooling Ratios
For reporting the structure preservation results described in Section 4.2, we considered multiple different proposals for what it means to preserve graph structure during pooling. In the end, we settled to compare the spectral distance between the symmetrically normalised graph Laplacians as an established measure of spectral property preservation, and investigated the relative difference in magnitude between the original graph $G$ and the pooled graph $G^{\prime}$ computed from diffusion distances. Specifically, the reported relative magnitude difference is calculated as
$\text{MagDiff}(G,G^{\prime})=\frac{|\text{Mag}(G)-\text{Mag}(G^{\prime})|}{\text{Mag(G)}}.$ (14)
For this experiment we further varied the pooling ratios across different pooling methods. However, some of the pooling layers considered in our study, specifically NDP, Graclus and NMF, were configured to always pool graphs to around half their size. To allow us to compare these methods across increasing pooling ratios, we choose to reapply these pooling operations repeatedly, which is why these three pooling methods are only evaluated at pooling ratios that are powers of 0.5.
C.5.2 Graph Classification Experiment
Our graph classification architecture follows the experimental setup described in Table 1 and is based the benchmark by Grattarola et al. [25]. Across our main classification experiment detailed in Table 1, different pooling layers are configured to reduce each input graph to around 50% of the number of nodes in the original graphs. Depending on the pooling method, this is chosen so each graph is reduced to 50% of its original size $k=\lfloor 0.5*N\rfloor$, (for NDP, Graclus, MagEdgePool, SpreadEdgePool, TopKPool, and SAGPool), or to 50% the average size of all graphs in the training dataset $k=\lfloor 0.5*\bar{N}\rfloor$ (for DiffPool, and MinCutPool). Interpreting the experimental results in Table 1 it is thus of interest that the sizes of the pooled graphs can vary across pooling layers, which might explain some of the difference in performance between the fixed-size methods DiffPool and MinCUT compared to more adaptive pooling methods.
Appendix D Extended Results
Finally, we summarise extended experimental results beyond the scope of our main paper.
D.1 Correlation between Magnitude and Spread
As stated in Section 2.2, the magnitude and spread of a metric space are closely related with magnitude giving an upper bound for spread when computed from the same positive definite metric space. Across our experiments on real graph datasets, we further find that this bound in practice can be very tight and magnitude and spread measure very related notions of effective size. More specifically, when computing both magnitude and spread from the diffusion distances detailed in Section 2.3, we observe that magnitude and spread almost coincide for all graphs from the NCI1, ENZYME or IMDB-Multi datasets as illustrated in Figure 5. In fact, magnitude and spread correlate almost perfectly across these three graph datasets (Pearson correlation $r^{2}\geq 0.99$). Further, we confirm that across these examples, magnitude is generally greater or equal to spread by a relatively low multiplicative factor close to 1. We therefore find empirical evidence for the fact that spread offers a valid and highly related alternative to magnitude in practice supporting our theoretical analysis of the relationship between spread and magnitude during pooling (Section B.3) as well as our observations on the similar performance of both MagEdgePool and SpreadEdgePool.
Figure 5: Comparison between magnitude and spread computed from diffusion distances for all graphs in three graph datasets, NCI1, ENZYMES and IMDB-Multi.
D.2 Evaluating Computational Efficiency
As detailed in Section B.2, the computational costs of our algorithm are determined by the costs of computing the edge scores used for pooling. To illustrate how this theoretical discussion translates into practice, we now investigate computational costs empirically in comparison to alternative pooling methods considered throughout this study.
Specifically, we first compare the runtimes of training the models used in Section 4.3 to compare the accuracy of different pooling methods across varying pooling ratios. Figure 6 then reports the mean runtime of training the GNN on one CV-fold for different choices of pooling ratios and pooling methods for the NCI1 dataset using either general convolutional layers or GIN layers. All models are trained as specified in Table 1 on a single GPU with 32GB memory.
Notably, we observe that MagEdgePool and SpreadEdgePool overall perform on par with alternative pooling methods in terms of runtimes. SpreadEdgePool has a consistent advantage over MagEdgePool due to the higher computational efficiency of computing spread rather than magnitude. Notice that for increasing pooling ratios, our algorithm re-computes the edge scores repeatedly, leading to a less pronounced decrease in computational costs than alternative methods. Nevertheless, we conclude that it is generally more efficient to apply SpreadEdgePool than to rely on trainable approaches, such as TopK and SAGPool for this dataset, indicating the computational benefit of non-trainable graph pooling operations. To generalise this runtime comparison to other datasets sizes and experimental setups, we note that our algorithm will scale with dataset size as described in Section B.2.
Figure 6: Runtime comparison for training the GNNs reported in the main results for NCI1 using different pooling layers. The left plot shows the mean time in seconds per run using general convolutional layers. The right plot shows the same result using GIN layers.
D.3 Node Feature Preservation and Expressivity
Graph pooling should not only preserve graph structure, but also preserve relevant node feature information during pooling. That is, pooling is frequently used after initial rounds of message passing and data representations learnt by previous layers should be respected and effectively encoded by the pooling procedure [7, 36]. One way of investigating node feature retention during pooling, is to evaluate how well a graph can be reconstructed from its pooled version. We follow the experimental setup by Grattarola et al. [25] to investigate. In particular, this experiment uses a model architecture, similar to the model proposed in Table 1, where each graph gets pooled to around 50% of nodes after the initial MLP and GNN layer. Then, the pooled graphs are up-scaled again by reversing the node selection step used by each pooling layer. From these unpooled graph representations, a further GNN and post-processing MLP layer are trained and the task is set to output the reconstructed node feature representation. This model is trained on each example graph using Adam to minimize the mean squared error (MSE) between the input and output node features with a learning rate of $0.0005$ and early stopping on the training loss with a patience of 1000 epochs and a tolerance of $10^{-6}$. Each experiment is repeated three times across different random seeds. See Grattarola et al. [25] for further explanations on the model architecture and experimental setup.
For this experiment, we expect SpreadEdgePool to perform comparably well as we specifically designed our pooling algorithm so that features are averaged during pooling. Further restricting the number of times a node can be merged effectively prevents the collapse of entire portions of the graph, which aids reconstruction. SpreadEdgePool pooling thus successfully encodes node information while allowing for a flexible choice of pooling ratio. This is confirmed by the results in Table 4, which highlight that SpreadEdgePool overall performs well at the features reconstruction task, especially for the sensor graph reaching low reconstruction errors. Alternative methods, in particular node drop approaches such as TopK and SAGPool, show notably worse feature preservation during this experiment indicating the benefits of more expressive pooling operations, such as SpreadEdgePool.
Note that the results in Table 4 capture one specific aspect of feature preservation, namely how well the features of specific example graphs can be reconstructed. This experiment does not assess the generalisation capability of pooling layers. Further, the experimental setup assumes that it is relevant to preserve all node features during pooling, which might not be realistic in practice, where the aim of pooling could be to solely encode task-relevant feature representations. Nevertheless, as discussed above, this extended experiment gives evidence to support that our proposed pooling algorithm, SpreadEdgePool, is capable of outputting expressive feature representations and aggregates node features in a faithful manner, which is likely one of the reasons for its high performance in graph classification tasks.
Table 4: Mean and standard deviation of the reconstruction MSE for reconstructing the original node positions from the pooled graph representations for different example graphs and pooling methods. Our proposed algorithm, SpreradEdgePool, does well at faithfully encoding the feature representations.
Ring Sensor barbell community erdosrenyi torus
SpreadEdge 5.47e-07 ± 2.63e-07 2.78e-05 ± 3.04e-07 3.42e-04 ± 1.42e-06 3.71e-03 ± 5.80e-05 6.49e-07 ± 3.33e-07 5.29e-07 ± 1.33e-07
NDP 3.08e-07 ± 3.57e-07 4.07e-05 ± 4.57e-06 4.54e-04 ± 2.01e-05 2.52e-01 ± 9.50e-06 1.46e-06 ± 1.18e-06 5.68e-07 ± 1.02e-07
Graclus 6.87e-04 ± 7.56e-07 2.67e-06 ± 2.31e-06 1.82e-03 ± 3.22e-07 2.42e+00 ± 2.11e-04 4.76e-02 ± 3.75e-07 7.10e-07 ± 8.60e-08
NMF 4.78e-07 ± 2.95e-07 1.96e-05 ± 1.39e-05 5.80e-04 ± 4.53e-07 6.06e-01 ± 1.43e-04 5.04e-07 ± 3.31e-07 2.52e-07 ± 2.93e-07
TopK 1.21e-01 ± 8.23e-03 5.83e-03 ± 2.16e-03 1.55e-02 ± 1.10e-02 6.03e+00 ± 2.21e+00 5.30e-03 ± 7.49e-03 1.72e-01 ± 8.51e-03
SAGPool 1.45e-01 ± 2.52e-02 2.01e-03 ± 2.73e-03 4.12e-02 ± 4.18e-02 4.76e+00 ± 1.57e+00 9.60e-05 ± 1.35e-04 1.88e-01 ± 4.63e-02
DiffPool 8.63e-06 ± 4.73e-06 3.50e-04 ± 8.32e-05 6.50e-04 ± 1.01e-06 2.14e-01 ± 2.84e-01 3.74e-04 ± 1.46e-04 5.17e-05 ± 8.90e-06
MinCut 2.55e-06 ± 2.68e-06 6.56e-06 ± 3.86e-06 2.35e-06 ± 1.50e-06 1.80e-04 ± 2.01e-04 1.44e-06 ± 5.24e-07 1.49e-06 ± 9.81e-07
D.4 Overview Experiment
To create the illustration in Figure 1 and visually compare the outputs of different pooling methods, we follow the experimental setup by Grattarola et al. [25] on understanding structure preservation in graph pooling layers. Specifically, we simulate a ring graph with 64 nodes, a barbell graph with 20 nodes on each side connected by 24 nodes in the middle, and a sensor graph with 64 nodes. All trainable pooling layers were then trained in a self-supervised manner to optimise the following spectral loss between the original graph $G$ and the pooled graph $G^{\prime}$:
$\mathcal{L}(G,G^{\prime})=\sum_{i=0}^{d}F^{\top}_{:,i}LF_{:,i}-F^{\prime\top}_{:,i}L^{\prime}F^{\prime}_{:,i}$ (15)
where $L,L^{\prime}$ are the corresponding Laplacian matrices. The features $F$ are taken to be the top 10 eigenvectors of $L$ concatenated with the coordinates of the nodes in $G$. $F^{\prime}$ is the reduced version of $F$ after pooling. Note that this type of spectral loss is one particular proposal on structure preservation and alternative objectives could be investigated.
Expanding on the qualitative comparison between the example graphs in Figure 1, Figure 7 shows quantitative structure preservation measures for all example graphs and pooling methods. Specifically, we summarise the spectral loss by Grattarola et al. [25], the spectral distance between the normalised graph Laplacians, and the magnitude difference between the pooled and original graphs as further detailed in Section 4.2.
Figure 7: Structure preservation measures for the examples in Figure 1. Pooling is repeated for three different random seeds and the annotations report the means and standard deviations of the structure preservation scores.
Figure 7 demonstrates that SpreadEdgePool and MagEdgePool do not only reach low magnitude differences across these three example graphs, they also show comparatively low spectral distances supporting our findings in Section 4.2. Further, we observe that methods that show worse visual preservation of graph structures, such as NMF, TopK, DiffPool, and MinCut, also reach consistently higher magnitude differences and spectral distances supporting our claim that these methods fail to faithfully preserve graph structures during pooling to varying extents.
D.5 Graph Classification
Table 5 further reports extended classification results on additional datasets extending on the results shown in Table 1. We chose not to report on these datasets in the main text because they showed fewer and less notable differences between pooling methods. Note that for the open graph benchmark MolHIV dataset, instead of using stratified cross-validation, we evaluate each model across predefined training, test and validation splits and evaluate their performance across 5 random seeds. Further, we report AUROC as the performance metric for MolHIV because is the suggested evaluation metric for this very imbalanced dataset. All other results are reported as in Table 1 via the mean and standard deviation of the test accuracy across 10-fold stratified cross-validation. In agreement with our main results, we observe that MagEdgePool and SpreadEdgePool constitute high performing general-purpose pooling methods that reach top performance across these extended datasets. In particular, for smaller biological datasets, such as MUTAG, BZR, or COX2, pooling via magnitude or spread notably improves on the GNN that uses no pooling layer, which indicates the beneficial effects of structure-aware pooling for graph learning.
Table 5: Classification performance of different pooling layers across multiple datasets. For each dataset, the best performing model is marked in bold and models that do not perform significantly different from the best performing model are coloured green.
Method MolHIV (AUROC) MUTAG COX2 BZR BZR_MD COX2_MD DHFR_MD ER_MD AIDS
No Pooling 74.4 ± 0.7 81.6 ± 6.3 83.8 ± 3.8 76.1 ± 9.1 73.7 ± 5.3 70.2 ± 8.1 71.3 ± 1.9 74.9 ± 0.8 99.0 ± 0.1
MagEdge 64.6 ± 8.7 84.0 ± 7.4 86.0 ± 7.3 88.1 ± 10.0 65.9 ± 2.2 76.5 ± 9.2 77.9 ± 9.4 79.9 ± 7.1 99.7 ± 0.1
SpreadEdge 70.1 ± 2.9 85.9 ± 7.4 85.1 ± 5.6 87.4 ± 9.5 69.7 ± 10.0 77.1 ± 9.0 74.6 ± 11.1 83.1 ± 4.0 99.7 ± 0.1
NDP 65.2 ± 7.3 91.6 ± 2.4 86.1 ± 7.1 86.3 ± 9.4 72.3 ± 11.3 73.9 ± 10.6 72.2 ± 11.4 84.2 ± 1.7 99.6 ± 0.1
Graclus 70.1 ± 5.3 87.4 ± 9.5 79.2 ± 12.5 80.1 ± 7.1 72.0 ± 9.6 75.4 ± 9.9 71.1 ± 12.3 81.7 ± 3.9 99.5 ± 0.1
NMF 73.2 ± 1.3 84.0 ± 9.4 83.8 ± 8.7 80.8 ± 8.5 69.0 ± 10.2 70.4 ± 6.6 70.8 ± 8.6 80.9 ± 2.0 97.0 ± 1.7
TopK 72.7 ± 1.8 81.9 ± 3.9 76.4 ± 7.2 75.8 ± 8.0 68.4 ± 8.5 65.9 ± 7.0 69.5 ± 3.3 74.0 ± 1.0 99.3 ± 0.1
SAGPool 74.3 ± 3.2 82.9 ± 2.3 76.8 ± 7.8 77.8 ± 5.0 66.8 ± 7.0 67.0 ± 6.4 70.5 ± 2.6 75.6 ± 1.2 99.0 ± 0.1
DiffPool 72.1 ± 1.0 83.3 ± 2.8 76.9 ± 6.5 76.8 ± 10.1 72.3 ± 3.0 69.2 ± 2.1 67.9 ± 2.6 73.4 ± 1.1 98.8 ± 0.2
MinCut 70.3 ± 3.0 80.6 ± 3.7 79.1 ± 4.4 70.8 ± 8.6 69.0 ± 5.0 70.0 ± 3.7 69.7 ± 1.9 73.3 ± 1.6 99.2 ± 0.1
D.6 Preserving Graph Structure
Extending on the results reported in Figure 3 and Section 4.2, we further report the distribution of structure preservation measures across all pooling ratios included in our main experiment. Specifically, Figure 8 shows line plots that summarise the mean magnitude difference relative to the original graph and the normalised spectral distance between all original and pooled graphs from the NCI1 dataset in the leftmost column. The remaining plots illustrate the quantiles of the same measures split up per pooling method. Overall, these individual plots support our assessment that MagEdgePool and SpreadEdgePool consistently reach low magnitude differences and comparably low spectral distances, with these trends being more pronounced in terms of the relative difference in magnitude after pooling.
Figure 8: Structure preservation for the NCI1 dataset across varying pooling ratios. Top row: The spectral distance between the normalised Laplacians of the original and the pooled graphs. Bottom row: The relative difference in magnitude, which summarises the proportional difference in structural diversity after pooling. Bold lines show the mean values of each score across graphs and thin lines the 10%, 25%, 75% and 90% quantiles.